The Raven Co. has just gone public. Under a firm commitment agreement, Raven received 18.50 for each of the 5 million shares sold. The initial offering price was $20.35 per share, and the stock rose to $26.25 per share in the first few minutes of trading. Raven paid $1,150,000 in direct legal and other costs, and $403,000 in indirect costs. What was the flotation cost as a percentage of funds raised?
Answer #1

Net amount raised =18.50*5000000 =92500000-1150000-403000 =90947000

share offered at price =20.35

amount received per share =18.50

underwriter spread =20.35-18.50=1.85 =1.85 per shares

total underwriter spread =per share spread*no of shares offered

1.85*5000000 = 9250000

total direct cost =9250000+1150000 =10400000

total direct flotation cost =10400000

indirect flotation cost =indirect cost given +price appreciation

total indirect cost =403000+(26.25-20.35)*5000000 =29903000

total flotation cost =direct cost +indirect cost =10400000+29903000=40303000

total flotation cost as a percentage of fund raised =(40303000/90947000)*100

=44.314%